# ICYM2I: The illusion of multimodal informativeness under missingness

### v2

@@ Acknowledgments
VJ, YSC, and SJ would like to acknowledge partial support from NIH R01MH137679. SJ would like to acknowledge partial support from the SNF Center for Precision Psychiatry and Mental Health at Columbia. Any opinions, findings, conclusions, or recommendations in this manuscript are those of the authors and do not reflect the views, policies, endorsements, expressed or implied, of any aforementioned funding agencies/institutions. The study was approved under IRB-AAAU7973 at Columbia University.
@@ Appendix B Partial Information Decomposition (PID)
Partial information decomposition (PID Williams and Beer [48]) consists in decomposing the total mutual information [60, 61] between a target variable and two input variables into information about the target variable that both input variables share (“Shared” information), only one input variable has (“Unique” information) and emerges from the interactions of both (“Complementary” information). Bertschinger et al. [47] introduces bounds for these, reiterated below. In this Appendix, we express these bounds as entropy. First, Table 5 summarizes the notations used in the literature and those used in our work.
@@
Bertschinger et al. [47] provides the following bounds on each of these quantities:
@@
In this context, Bertschinger et al. [47] demonstrates that solving the optimization for $q\in{\Delta_{\Omega^{\text{}}}}$ that satisfies one of the four conditions above or the bound for conditional entropy, $\widetilde{H}$ formalized in (5), is sufficient to obtain all the quantities of interest.
$\widetilde{H}(Y|X_{1},X_{2})$ $=\max_{q\in{\Delta_{\Omega^{\text{}}}}}H_{q}(Y|X_{1},X_{2})$ (5)
Formulating PID quantities in terms of entropy
Remember that entropy $H(\cdot)$ is defined for general distributions of $X$ and $Y$ as follows:
@@
where the first equation comes from the chain rule of mutual information [76].
@@
Appendix C ICYM2I-PID
@@
Training [Uncaptioned image]
Current
practice
IPW-adjusted
evaluation alone
IPW-adjusted
training alone
ICYM2I  (IPW-adjusted
training and evaluation)
Let $\text{IPW}_{q}(p)$ denote IPW reweighting to correct $p\mapsto q$ using samples from $p$. We use the following projection set to operationalize PID-bound estimation:
$\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$
$\approx$ $\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(X_{i}=x_{i},Y=y)={p_{\Omega^{\text{}}}}_{\phi}(y,x_{i})\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
$=$ $\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(X_{i}=x_{i},Y=y)=\text{IPW}_{{p_{\Omega^{\text{}}}}_{\phi}}({p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(y,x_{i}))\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
Note that this differs from prior works [46] that have used:
$\Delta_{\Omega^{\text{}}}\approx\{q\propto\exp(f_{1}(x_{1})\cdot f_{2}(x_{2})):q(x_{i},y)={\Omega_{\text{obs}}^{\text{}}}_{\phi}(x_{i},y)\,\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\}$
To further improve the robustness of $p_{\Omega_{\text{obs}}^{\text{}}}(y,x_{i})\approx{p_{\Omega_{\text{obs}}^{\text{}}}}_{\phi}(y,x_{i})$, we optionally leverage both the unimodal and multimodal estimations, i.e., the unimodal and multimodal prediction models to enforce:
$\sum_{x_{\neg i}}f_{\phi}(y\mid x_{1},x_{2})p_{\Omega_{\text{obs}}^{\text{}}}(x_{1},x_{2})=\hat{f}_{\phi}(y\mid x_{i})p_{\Omega_{\text{obs}}^{\text{}}}(x_{i}),$
where $\hat{f}$ is obtained by adding the smallest amount of noise to $f_{\phi}(y\mid x_{i})\approx p_{\Omega_{\text{obs}}^{\text{}}}(y\mid x_{i})$ to ensure that the joint probabilities from the parametrized unimodal and multimodal models match exactly. This optimization did not improve PID values’ recovery.
We summarize our proposed auto-differentiable PID algorithm with IPW-based correction in Algorithm 1.
@@
1:# Adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$.
@@
3: Estimate corrected unimodal and multimodal models by training with weighting IPW-loss: $f_{\phi}(y\mid x_{i})\approx p_{\Omega^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$, and $f_{\phi}(y\mid x_{1},x_{2})\approx p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})$.
4: # Drop subscript $\phi$ from $f_{\phi}$ for clarity.
5: Initialize parameterizations $\theta$ for $q$: $f_{i}(y\mid x_{i}),\,\forall i\in\{1,2\}$.
@@
7:while not converged do
8:  for samples in batch do
9:    # Project onto $\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$.
10:   $q_{\theta}(y\mid x_{1},x_{2})\leftarrow\texttt{SINKHORN-KNOPP}(q_{\theta}(y\mid x_{1},x_{2}),\{p(y,x_{i})\}_{i=1}^{2})$.
11:   Estimate the loss $I_{q}(Y:(X_{1},X_{2}))$ as a batch sample mean.
12:   $\theta\leftarrow\theta-\nabla_{\theta}I_{q}(Y:(X_{1},X_{2}))$.
13:  end for
14:end while# Estimate mutual information under $p_{\Omega^{\text{}}}$.
15: Estimate $I_{\Omega^{\text{}}}(Y:(X_{1},X_{2})),\,\text{and }I_{\Omega^{\text{}}}(Y:X_{i}),\,\forall i\in\{1,2\}$ using adjustment in Appendix A.
@@
17:return $PID(\Omega^{\text{}})$
The traditional SINKHORN-KNOPP algorithm updates a matrix to enforce its marginals to be unit vectors. In our work, we adapt the algorithm to enforce the marginals to match $p_{\Omega}^{\text{}}$-marginals, ensuring that $q_{\theta}(\cdot)\in\Delta_{\Omega}^{\text{}}$. To ensure proper gradient propagation and reduce memory use, we use the unrolled SINKHORN-KNOPP [77, 78] algorithm. In the following, we use subscripts $q_{x_{1},x_{2}}$ to denote $q_{\theta}(y,x_{1},x_{2})$ and $p_{x_{i}}$ to denote $p_{\phi}(y,x_{i})$. The algorithm is detailed below:
@@
5:  if $\left|\frac{q_{x_{1}}-p_{x_{1}}}{p_{x_{1}}}\right|\leq\texttt{atol}$ $\left|\frac{q_{x_{2}}-p_{x_{2}}}{p_{x_{2}}}\right|\leq\texttt{atol}$ then
@@ Appendix E Synthetic data results
Data generation. Our work builds on the example introduced in [46], in which we enforce additional missingness. Three latent variables ($z_{1}$, $z_{2}$, and $z_{c}$) are drawn from multi-dimensional clustered data; the observed covariates are a concatenation of $z_{c}$ and one of the other latent variables, as illustrated in Figure 3. Then, the outcome $Y$ is generated as $Y=\sigma(p_{1}\mathbb{E}(z_{1})+p_{2}\mathbb{E}(z_{2})+(1-p_{1}-p_{2})\mathbb{E}(z_{c}))$, with the proportion $p_{i}\in[0,1]$ such that $p_{1}+p_{2}\leq 1$. We simulate datasets with varying values of $p_{1}$ and $p_{2}$. Then, we enforce a $50\%$ MAR missingness pattern in $X_{2}$ by modeling the probability of missingness. We do this by clustering $X_{1}$ into 100 groups using Kmeans. Then, the probability of missingness is generated using a random forest that regresses $X_{1}$ to predict $c_{j}\cdot{Y}$.
@@
Empirical setting. Data were split into three: 80% for training, 10% for validation, and the rest for testing. We consider neural networks with 2 hidden layers with 32 nodes, trained using an Adam optimizer [79] with a learning rate of 0.001 over 100 epochs. Our evaluation relies on discriminative performance measured through AUROC.
Estimating predictive performance under $\Omega_{\text{obs}}^{\text{}}$. Table 9 presents the estimated performance obtained under different corrections. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the performance one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. Note that in this setting, we rely on the true IPW correction that one would obtain with a properly specified model, as the MAR setting is met.
Table 9: Comparison between estimated AUC performance under the different training and evaluation corrections and oracle performance on $\Omega^{\text{}}$. $\epsilon$ denotes the RMSE between estimated and oracle PIDs.
Evaluation
Observed Underlying
$\Omega_{\text{obs}}^{\text{}}$ $\Omega^{\text{}}$
[Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@ F.1 UR-FUNNY
We illustrate the impact of missingness on estimating the informativeness of different modalities on real-world data with UR-FUNNY [64], a multimodal dataset for humor detection from human speech used in affective computing. The dataset comprises text, audio, and visual modalities from 10 - 20 second videos sourced from TED talks, and the task is to detect whether a punchline would trigger a laugh. Labels were generated using the markup “(Laughter)” [80] from the transcript.
Dataset. The processed dataset from MultiBench [81] is a modality-complete dataset with 10,166 samples of paired audio, text, and vision embeddings. Audio embeddings were generated with COVAREP [82], text with Glove [83], and visual features through the Facet [84] library and OpenFace [85], and aligned using the Penn Phonetics Lab Forced Aligner (P2FA) [84].
@@
This figure shows that the proposed strategy is consistently closer to the Oracle’s PID values. This demonstrates that under Assumption C, the proposed correction yields better estimates of each modality’s informativeness – specifically, the audio-visual modality (Unique 1) carries more information.
@@ F.2 Hateful Memes
We run experiments using the dataset from the Hateful Memes Challenge [65], which investigates text-image multimodal reasoning in the context of hate speech detection in online memes. The dataset comprises text-image pairs with an associated label indicating hate speech.
Dataset. We utilize the Kaggle version of the Facebook Hateful Memes dataset, as referenced in the Holistic Evaluation of Multimodal Foundation Models (HEMM) [86] repository. Our analysis focuses on the 9,000 samples with associated labels. For each sample, embeddings were extracted for both modalities using a ResNet-50 [87] for images and a BERT-base-uncased [88] model for text. The proposed ResNet-50 was pretrained on ImageNet [89] with the final layer replaced to extract 2048-dimensional feature vectors, and BERT-base-uncased [88] extracts embeddings of dimension 784 from the penultimate layer.
@@ Appendix G Structural heart disease data processing
All electrocardiograms were 10-second, standard 12-lead ECG signals collected at abstracted to 250 Hz, which we resampled to 500 Hz, and standard normalized by channel to match the inputs for ECG-FM [71]. We used the version of ECG-FM with weights pretrained on MIMIC-IV [90, 91] and PhysioNet 2021 [92, 93]. We averaged the outputted feature embeddings along the temporal dimension and flattened them to produce vectors of length 768.
The chest radiographs used in our study were all postero-anterior (PA) view CXRs. We extracted pixel values from the DICOM files as grayscale images, center-cropped each image along the shorter dimension, applied contrast-limited adaptive histogram equalization (CLAHE) [94] with a clip limit of 0.2, and resized each image to $1284\times{}1284$ pixels. All outputted embeddings were flattened to 4098-dimensional vectors.

### v3

@@ Acknowledgments
VJ and SJ acknowledge partial support from NIH R01MH137679. YSC and SJ acknowledge partial support from Google Research Scholar Awards. SJ acknowledges partial support from the SNF Center for Precision Psychiatry and Mental Health at Columbia. Any opinions, findings, conclusions, or recommendations in this manuscript are those of the authors and do not reflect the views, policies, endorsements, expressed or implied, of any aforementioned funding agencies/institutions.
@@ Appendix B Partial Information Decomposition (PID)
Partial information decomposition (PID Williams and Beer (2010)) consists in decomposing the total mutual information (McGill, 1954; Te Sun, 1980) between a target variable and two input variables into information about the target variable that both input variables share (“Shared” information), only one input variable has (“Unique” information) and emerges from the interactions of both (“Complementary” information). Bertschinger et al. (2014) introduces bounds for these, reiterated below. In this Appendix, we express these bounds as entropy. First, Table 5 summarizes the notations used in the literature and those used in our work.
@@
Let $\Delta$ be the space of all distributions over $(X_{1},X_{2},Y)$ and let $\Omega^{\text{}}$ denote the true data distribution (without missingness) and define $\Delta_{\Omega^{\text{}}}:=\big\{q\in{\Delta}:q(X_{i}=x_{i},Y=y)=p_{\Omega}^{\text{}}(X_{i}=x_{i},Y=y)\text{ }\forall x_{i}\in{\mathcal{X}_{i}},y\in{\mathcal{Y}},i\in\{1,2\}\big\}$. That is, $\Delta_{\Omega^{\text{}}}$ is the set of all distributions over $(X_{1},X_{2},Y)$ such that the two-way joints between $X_{i}$ and $Y$ match the true data-generating distribution. Equipped with this set, Bertschinger et al. (2014) provides the following bounds $\widetilde{SI}$, $\widetilde{UI}$, and $\widetilde{CI}$ on the analogous quantities:
@@
In this context, Bertschinger et al. (2014) demonstrates that solving the optimization for $q\in{\Delta_{\Omega^{\text{}}}}$ that satisfies one of the four conditions above is sufficient to obtain all the quantities of interest. Importantly, these bounds are tight if there exists a $q_{0}\in\Delta_{\Omega^{\text{}}}$ such that $\widetilde{CI}_{q_{0}}(Y:X_{1};X_{2})=0$. Bertschinger et al. (2014) further shows that under common (but unverifiable) assumptions on the data-generating process, the inequalities are tight for all $q\in\Delta_{\Omega^{\text{}}}$. This results in a compelling argument for relying on these entities, as it suggests that it is not possible to decide whether complementary information exists when only the marginals $(Y,X_{1})$ and $(Y,X_{2})$ are known.
Formulating PID quantities in terms of entropy.
For stability, we propose to formalize the previous bound in terms of entropy, $H(\cdot)$, defined for general distributions of $X$ and $Y$ as follows:
@@
where the first equation comes from the chain rule of mutual information (Wyner, 1978).
@@
Appendix C ICYM2I
C.1 ICYM2I-Learn
Table 6 summarizes the proposed approach to estimate performance using $\Omega_{\text{obs}}^{\text{}}$. Critically, one must correct both training and evaluation to obtain the performance on $\Omega^{\text{}}$.
@@ C.1 ICYM2I-Learn
Training [Uncaptioned image] Current practice IPW-adjusted evaluation alone
[Uncaptioned image] IPW-adjusted training alone ICYM2I  (IPW-adjusted training and evaluation)
C.2 ICYM2I-PID
The following describes the algorithm to estimate PID under missingness.
@@ C.2 ICYM2I-PID
1:# Step 1: Adjust $\Omega_{\text{obs}}^{\text{}}\mapsto\Omega^{\text{}}$.
@@
3:# Step 2: Train corrected unimodal and multimodal models.
4: Training each model with weighting IPW-loss: $f(y\mid x_{i})\approx p_{\Omega^{\text{}}}(y\mid x_{i}),\,\forall i\in\{1,2\}$, and $f(y\mid x_{1},x_{2})\approx p_{\Omega^{\text{}}}(y\mid x_{1},x_{2})$.
5:# Step 3: PID optimization.
6: Initialize parameterizations $\theta$ for $q$: $f_{i}(y\mid x_{i}),\,\forall i\in\{1,2\}$.
@@
8:while not converged do
9:  for samples in batch do
10:    # Ensure $q\in\Delta_{\Omega^{\text{}}}^{\textsc{ICYM${}^{2}$I}}$ by projection.
11:   $q_{\theta}(y\mid x_{1},x_{2})\leftarrow\texttt{SINKHORN-KNOPP}(q_{\theta}(y\mid x_{1},x_{2}),\{p_{\Omega}^{\text{}}(y,x_{i})\}_{i=1}^{2})$.
12:   Estimate the loss $I_{q}(Y:(X_{1},X_{2}))$ as a batch sample mean.
13:   $\theta\leftarrow\theta-\nabla_{\theta}I_{q}(Y:(X_{1},X_{2}))$.
14:  end for
15:end while# Step 4: Estimate mutual information under $p_{\Omega^{\text{}}}$.
16: Estimate $I_{\Omega^{\text{}}}(Y:(X_{1},X_{2})),\,\text{and }I_{\Omega^{\text{}}}(Y:X_{i}),\,\forall i\in\{1,2\}$ using adjustment in Appendix A.
@@
18:return $PID(\Omega^{\text{}})$
The traditional SINKHORN-KNOPP algorithm updates a matrix to enforce its marginals to be unit vectors. In our work, we adapt the algorithm to enforce the marginals to match $p_{\Omega}^{\text{}}$-marginals, ensuring that $q_{\theta}(\cdot)\in\Delta_{\Omega}^{\text{}}$. To ensure proper gradient propagation and reduce memory use, we use the unrolled SINKHORN-KNOPP (Sinkhorn and Knopp, 1967; Cuturi, 2013) algorithm. In the following, we use subscripts $q_{x_{1},x_{2}}$ to denote $q_{\theta}(y,x_{1},x_{2})$ and $p_{x_{i}}$ to denote $p_{\phi}(y,x_{i})$. The algorithm is detailed below:
@@
5:  if $\left|\frac{q_{x_{1}}-p_{x_{1}}}{p_{x_{1}}}\right|\leq\texttt{atol}$ and $\left|\frac{q_{x_{2}}-p_{x_{2}}}{p_{x_{2}}}\right|\leq\texttt{atol}$ then
@@ Appendix E Synthetic data results
Data generation. Our work builds on the example introduced in (Liang et al., 2024a), in which we enforce additional missingness. Three latent variables ($z_{1}$, $z_{2}$, and $z_{c}$) are drawn from multi-dimensional clustered data; the observed covariates are a concatenation of $z_{c}$ and one of the other latent variables, as illustrated in Figure 3. Then, the outcome $Y$ is generated as $Y=\sigma(p_{1}\mathbb{E}(z_{1})+p_{2}\mathbb{E}(z_{2})+(1-p_{1}-p_{2})\mathbb{E}(z_{c}))$, with the proportion $p_{i}\in[0,1]$ such that $p_{1}+p_{2}\leq 1$. We simulate datasets with varying values of $p_{1}$ and $p_{2}$. Then, we enforce a $50\%$ MAR missingness pattern in $X_{2}$ by modeling the probability of missingness. We do this by clustering $X_{1}$ into 100 groups using Kmeans. Then, the probability of missingness is generated using a random forest that regresses $X_{1}$ to predict $c_{j}\cdot{Y}$.
@@
Empirical setting. Data were split into three: 80% for training, 10% for validation, and the rest for testing. We consider neural networks with 2 hidden layers with 32 nodes, trained using an Adam optimizer (Kingma and Ba, 2014) with a learning rate of 0.001 over 100 epochs. Our evaluation relies on discriminative performance measured through AUROC.
Estimating predictive performance under $\Omega_{\text{obs}}^{\text{}}$. Table 9 presents the PID obtained under different corrections. These results underline the importance of correcting both training and evaluation, as proposed in ICYM2I, to best align with the PID one would obtain on $\Omega^{\text{}}$, as shown by the smallest Root Mean Squared Error (RMSE) observed when both corrections are applied. Note that in this setting, we rely on the true IPW correction that one would obtain with a properly specified model, as the MAR setting is met.
Observed ($\Omega_{\text{obs}}^{\text{}}$) Underlying ($\Omega^{\text{}}$)
$\Omega^{\text{}}$ [Uncaptioned image] [Uncaptioned image]
@@ F.1 UR-FUNNY
We illustrate the impact of missingness on estimating the informativeness of different modalities on real-world data with UR-FUNNY (Hasan et al., 2019), a multimodal dataset for humor detection from human speech used in affective computing. The dataset comprises text, audio, and visual modalities from 10 - 20 second videos sourced from TED talks, and the task is to detect whether a punchline would trigger a laugh. Labels were generated using the markup “(Laughter)” (Chen and Lee, 2017) from the transcript.
Dataset. The processed dataset from MultiBench (Liang et al., 2021) is a modality-complete dataset with 10,166 samples of paired audio, text, and vision embeddings. Audio embeddings were generated with COVAREP (Degottex et al., 2014), text with Glove (Pennington et al., 2014), and visual features through the Facet (Yuan et al., 2008) library and OpenFace (Baltrušaitis et al., 2016), and aligned using the Penn Phonetics Lab Forced Aligner (P2FA) (Yuan et al., 2008).
F.1.1 Missing At Random (MAR)
@@
This figure shows that the proposed strategy is consistently closer to the Oracle’s PID values. This demonstrates that under Assumption A, the proposed correction yields better estimates of each modality’s informativeness – specifically, the audio-visual modality (Unique 1) carries more information.
F.1.2 Missing Not At Random (MNAR)
Enforcing missingness. A central assumption of our method is that missingness is MAR. We propose to analyze the impact of violations of this assumption, specifically the presence of MNAR patterns, on the quality of estimates obtained using our correction. To this end, we simulate audio and visual missingness as a function of the modality itself. Similarly to the previous analysis, we vary missingness from 30% to 70%. To estimate propensity in this setting, we rely on a logistic regression model based on the fully observed modality.
Results. Figure 5 shows the PID values obtained under the three previously described strategies. Critically, the proposed correction leads to performance similar to the model without correction, as the missingness probabilities cannot be estimated from the observed modality. This example illustrates the importance of assessing the plausibility of Assumption B in real-world settings, as no theoretical guarantees hold in such settings.
@@ F.2 Hateful Memes
We run experiments using the dataset from the Hateful Memes Challenge (Kiela et al., 2020), which investigates text-image multimodal reasoning in the context of hate speech detection in online memes. The dataset comprises text-image pairs with an associated label indicating hate speech.
Dataset. We utilize the Kaggle version of the Facebook Hateful Memes dataset, as referenced in the Holistic Evaluation of Multimodal Foundation Models (HEMM) (Liang et al., 2024b) repository. Our analysis focuses on the 9,000 samples with associated labels. For each sample, embeddings were extracted for both modalities using a ResNet-50 (He et al., 2016) for images and a BERT-base-uncased (Devlin et al., 2019) model for text. The proposed ResNet-50 was pretrained on ImageNet (Deng et al., 2009) with the final layer replaced to extract 2048-dimensional feature vectors, and BERT-base-uncased (Devlin et al., 2019) extracts embeddings of dimension 784 from the penultimate layer.
@@ Appendix G Structural heart disease data processing
G.1 Embedding generation
We generate embeddings using modality-specific foundation models–ECG embeddings are generated using ECG-FM (McKeen et al., 2024) and CXR embeddings with ELIXR-C (Xu et al., 2023).
All electrocardiograms were 10-second, standard 12-lead ECG signals collected and abstracted to 250 Hz, which we resampled to 500 Hz, and standard normalized by channel to match the inputs for ECG-FM (McKeen et al., 2024). We used the version of ECG-FM with weights pretrained on MIMIC-IV (Johnson et al., 2023; Goldberger et al., 2000) and PhysioNet 2021 (Reyna et al., 2021; 2022). We averaged the output feature embeddings along the temporal dimension and flattened them to produce vectors of length 768.
The chest radiographs used in our study were all postero-anterior (PA) view CXRs. We extracted pixel values from the DICOM files as grayscale images, center-cropped each image along the shorter dimension, applied contrast-limited adaptive histogram equalization (CLAHE) (Pizer et al., 1987) with a clip limit of 0.2, and resized each image to $1284\times{}1284$ pixels. All outputted embeddings were flattened to 4098-dimensional vectors.
G.2 IPW Correction
To address missingness in the observed CXRs, we apply the proposed propensity-based correction. The propensity scores are obtained from a logistic regression model using the ECG embedding, along with sex and age as predictors, serving as proxies for the socio-medical factors that influence whether a CXR is collected. Controlling for these covariates aims to render the MAR assumption more plausible. In practice, all relevant covariates, even outside of modalities being modeled, can be used for the correction.